OpenStudy (anonymous):

yooo can someone help me with evaluating the expression by sketching a triangle, sin(2cos -1 3/5)

OpenStudy (anonymous):

is this the expression: \(\large sin(2cos^{-1}(\frac{3}{5})) \) ???

OpenStudy (anonymous):

if yes, then let \(\large y=cos^{-1}\frac{3}{5} \) so cosy = 3/5. draw a right triangle depicting that trig relation: